OpenStudy (anonymous):

Jane has a 5 × 7-inch photograph that she wants enlarged to a 10 × 14-inch photograph. What is the scale factor from the original photograph to the enlarged photograph?

OpenStudy (triciaal):

how do you get from 5 to 10 or from 7 to 14? re -worded 10/5 = ?

OpenStudy (anonymous):

5*2=10 10/5=2